Session: May 2002 (session 2)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Carlo Migoni at Wed May 29 17:36:10 CEST 2002) Success

Noto (P. Leto at Thu May 30 13:30:36 CEST 2002) Success
No known problems

Jodrell_Bank (Alastair Gunn at Thu May 30 16:26:20 CEST 2002) Minor failures
VC03 unlocked and swapped for VC14 at about 09:06 UT. No other known problems.

Cambridge (Alastair Gunn at Thu May 30 16:27:39 CEST 2002) Success
No known problems.

Torun (K.Borkowski at Fri May 31 06:55:42 CEST 2002) Minor failures
The first tape was badly packing and apparently this led the recorder to a stop at 8:39 (resumed at 9:08). 2nd tape better packing yet not very nice. BBC6 unstable.

Effelsberg (A. Kraus at Fri May 31 12:37:48 CEST 2002) Minor failures
By accident, the cal was switched on until UT0800.

Westerbork (Tony Foley at Mon Jun 3 09:45:26 CEST 2002) Success
No known problems

Onsala (John Conway at Mon Jun 3 14:25:53 CEST 2002) Minor failures
No known problems